Latin Corporation’s breakeven point in revenues is $1,000,000 and fixed costs is


$400,000


Required:


1) Compute the contribution margin ratio


2) Compute the selling price if variable costs are $12 per unit


3) Suppose 80,000 units are sold, compute the margin of safety

1
Expert's answer
2022-01-18T13:02:15-0500

1) Contribution margin ratio:

"CMR=\frac{400000}{1000000}=0.40"


2) Selling price:

If variable costs are $12 per unit then selling price is:

"Q=\frac{VC}{AVC}=\frac{600000}{12}=50000 units"



"P=\frac{400000}{50000}+12=20"



3) Margin of safety:

If 80,000 units are sold then the margin of safety is:

"MOS=\frac{80000-50000}{80000}=0.37"